40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 15 12:25:38 2011, kiwamu, Update, Green Locking, Y arm ALS : beat-note free run fluctuation noise_budget.png
    Reply  Wed Nov 16 17:29:46 2011, kiwamu, Update, Green Locking, Some updates on the Y end green PDH Yarm_fluctuation.png
       Reply  Thu Nov 17 17:21:22 2011, kiwamu, Update, Green Locking, Y end green PDH servo : it's okay servoTF.pngservofilter.pngerr_suppression.png
          Reply  Fri Nov 18 11:28:39 2011, rana, Update, Green Locking, Y end green PDH servo : it's okay 
             Reply  Fri Nov 18 12:11:24 2011, Zach, Update, Green Locking, Y end green PDH servo : it's okay 
    Reply  Thu Nov 17 18:20:26 2011, kiwamu, Update, Green Locking, Y arm ALS : 1st trial of noise budget Yarm_ALS_2011Nov16.png
       Reply  Mon Nov 21 16:08:04 2011, kiwamu, Update, Green Locking, 2nd trial of Y arm ALS noise budget : broad band noise gone Yarm_ALS_2011Nov19.png
          Reply  Tue Nov 22 15:18:18 2011, kiwamu, Update, Green Locking, broad band noise depends on the gain of Y green PDH. and comaprator broken 
Message ID: 5914     Entry time: Wed Nov 16 17:29:46 2011     In reply to: 5894     Reply to this: 5929
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Some updates on the Y end green PDH 
Quote from #5894

 (Things to be done)

   [DONE]   1.1 Measurement of the arm fluctuation => to allow re-designing the servo shape
   [DONE]   1.2 temporary SR560 servo
   [ONGOING]1.3 Sanity checks on the modulation depth, reflectivity, PD dark noise and etc.,
   [DONE]  1.4 Make the servo more robust
   [DONE]  1.5 Some modifications on the medm screens
   [NOTYET]   1.6 Activation of the temperature feedback through the realtime digital control

Some updates on the Y end green PDH lock

(Measurement of the Y arm fluctuation)

In order to design the PDH box's servo shape we wanted to measure the Y arm fluctuation.
Here is the spectrum taken by looking at the control signal before the laser PZT.
 
 Yarm_fluctuation.png
 The scale of the Y axis is calibrated by using the PZT response of 5 MHz/V.
Above 10 Hz the spectrum shows 1/f noise which I believe the laser frequency noise.
 

(Temporary servo setup)

 We have found that the servo shape was not enough (#5890) to well-suppress the fluctuation shown above.
 Since the Newfocus fast servo box only makes 1/f shape, the error signal wasn't suppressed within the linear range.
So I have added an SR560 in the other input of the Newfocus servo box to make the filter shape 1/f^2.
Then the lock became more solid and the reflected DC light in time series is now much flat if the alignments are good.
I will post the servo shape and diagram later.

(Sanity checks)

 I looked at the reflected DC light when the laser was kept locked.
The reflectivity of the Y arm cavity went down to about 30% and this is good because it is supposed to be 27.5% when it is locked according the spec.
This means the mode-matching is not so bad.
ELOG V3.1.3-